Systems Staff Meeting Notes 10/07/02

Systems Staff Meeting Notes
October 7, 2002

Present: Rich Bennett, Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
Announcements
  1. Bobbie’s daughter won second place in barrel racing (horses).
  2. October 31 – Systems Halloween –freelance costume dress-up.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. 23 new machines for the Staff Training classroom in MSL are on order. They will small and fast – 2.4 Ghz, 80G HD, 1 G memory – with flat screen panels. Date of arrival is unknown.
  2. Staff Training classroom – they are trying to get the furniture done. Need to get plans for the physical layout from Facilities, then contact Judy (Digital crew). We’ll need 23 drops on the 3rd floor of MSL – also switches. It will have 22 workstations, 2 tables. Duplex – 5 places 3’ width 2’ depth. Tables/cable management – farthest 6’ – run to wall jack.
  3. Maps server for DLC to be purchased. ARC info – will be a new baby Dell, to be housed in Systems. Metadata will be manipulated commercially.
  4. Exchange server outage over the weekend – scratch disk space on the C drive ran out of space. Log entries show that the biggest single culprit was the virus scanner. Had a lot of Bug Bear and Opera F viruses. Both were being caught at the server or the workstation levels. There was a communication problem with informing Systems – no one called Bill at home. Supposedly someone was paged but did not answer. There were several email messages attempting to inform SYSHELP, but obviously those were not retrievable.
  5. Temperature alarm problems in the tape vault – the alarm kept going off on Friday and over the weekend. We keep resetting it, but probably need to get Facilities over to look at it.
  6. Xerox machine vandalism – the video servers in Facilities mysteriously turned off right before the incident in MSL occurred. The West videos should have been recording. As of Friday, both building videos were working. BellStor needed defragging – usually takes 24 hours, so Will just did a quick format.
Networking
  1. User permissions – Wednesday deadline for NT2R – need to check all old files on NT2 to doublecheck that they’ve been moved to NT2R.
  2. Two auxiliary servers for Facilities are in place – have IPs – need to plug up and install the database. Rob Roberts is out this week. Bill would like it to be in place and ready to go – setting the screensaver to the 3 minute lock up.
  3. Active Directory – 3 1/2 weeks until it is supposed to go campus-wide.
Software Projects Status
  1. PAULA – still dies. Cynthia is checking into it. Seems to happen on Bill’s machine especially.
  2. Disappearing log-off button on public workstations is still occurring. Could be related to the PAULA problem. Checked on the Robena tip, but it did not help resolve the matter.
  3. Red Alert button on PAULA – is done, but still testing.
  4. SysRequest – ready to go. Debra will be entering the data; it won’t work with the students. At this time, the next person in line for approval will need to be informed to check the SysRequest Web page.
  5. Hilton attended a non-technical meeting on the campus-wide SSN removal. Basically it was an overview, but the kickoff month was announced: January 2003. The Registar’s Office has pleaded that the change not take place when the students first come back after the holidays. For that reason, it has been moved to MLK weekend. Rich notes that this new policy was announced on Channel 20 news. The report said that students’ SSN would be collected, but not used.
  6. DLC – Ying had another mod to do for the Tracking database. She had a meeting with Erich. He had 3 requests that were logically impossible. She tried to get him to understand and they finally came to a solution.
  7. DLC – OPS reporting – meeting with Ying, Bill and Winston. Erich wants to measure how much time his students spend on a project. He feels that their productivity is too low. Bill says that Resource Services has a database to record time spent on projects. Erich has a prototype to show on Wednesday. Hilton may have worked on it in the past
  8. DLC – dates disappearing in Tracking Database – Ying had not seen any inconsistencies in the database until the student made changes. In the future, if we find that anyone has manipulated one of our databases, Bill needs to be informed.
  9. Special OPS Report – Cynthia and Winston demoed it – Barbara said it was OK – Steve Shorb had one minor change, then it will be ready to go live. Grace didn’t like it and Evelyn seemed a bit blown away by it – the different categories and divisions. The two versions now mesh on Divisions. This is part of what Erich wants. This is the expense part – the other is to measure productivity. Need to have a live copy of it in hand when we talk to Erich.
  10. Caribbean Newspaper mods are done.
  11. Hilton is working on a program to convert the old IDs (SSN) to the new UF IDs, then delete the SSNs.
Web Projects
  1. Suzy checked the Stat server – it stopped over the weekend. We’re using the most up-to-date version.
  2. We will stop the normal backups for it to see if it works. Suzy says not to backup NT8R for one week to see if Stat function restores.


Back to the Systems Homepage

Last updated October 14, 2002
by Debra Harris
debharr@mail.uflib.ufl.edu